one gram of ice at 0°C is mixed with one gram of water at 100°C the resulting temperature will be
1)5°C
2)0°C
3)10°C
4)100°C
1
Expert's answer
2019-05-03T09:00:50-0400

"Q_1=Q_2+Q_3"

"Q_1" - amount of heat transferred from the water cooling

"Q_2" - amount of heat transferred for the ice melting

"Q_3" - amount of heat transferred for the melted ice heating

"Q_1=m_w c_w (T_{w} - T)"

"Q_2=m_i c_i"

"Q_3=m_i c_w (T-T_i)"

"m_w" - mass of the water

"m_i" - mass of the ice

"T_w" - initial temperature of the water

"T_i" ​ - initial temperature of the ice

"T" - resulting temperature of the mixture

"c_w" - specific heat of the water, "c_w \approx 4.186 \; J/(g \cdot \degree C)"

"c_i" - specific latent heat of fusion of the ice, "c_i \approx 334 \; J/g"

"1 \cdot 4.186 \cdot (100-T)=1 \cdot 334 + 1 \cdot 4.186 \cdot (T-0)"

"418.6-4.186T=334 + 4.186T"

"8.372T=84.6"

"T= \frac{84.6}{8.372}"

"T \approx 10 \degree C"


Answer: "10 \degree C".